OpenStudy (anonymous):

Trigonometry help Multiply and Simplify sinx cosx (cscx + tanx)= Please show me how to solve this, I've been struggling with this.

OpenStudy (imstuck):

I can help!

OpenStudy (anonymous):

Awesome! Please do!

OpenStudy (imstuck):

First tell me what csc is the cofunction for.

OpenStudy (anonymous):

like 1\sin x?

OpenStudy (imstuck):

Then do the same with tangent. Do you know what I mean by that?

OpenStudy (anonymous):

and tan is sinx\cosx?

OpenStudy (imstuck):

Yes to both!

OpenStudy (imstuck):

so let's rewrite our expression using those identities instead:

OpenStudy (imstuck):

\[\sin x \cos x (\frac{ 1 }{ \sin x }+\frac{ \sin x }{ \cos x })\]

OpenStudy (anonymous):

I have gotten as far as: sinx cosx \[\frac{ 1 }{ sinx }\] + sinx cosx \[\frac{ sinx }{ cosx }\]

OpenStudy (anonymous):

i cant get it to type on one line but that

OpenStudy (imstuck):

Now you have to distribute the sin(x)cos(x) into the parenthesis like this:

OpenStudy (imstuck):

\[\frac{ \sin(x)\cos(x) }{ \sin(x) }+\frac{ \sin ^{2}(x)\cos(x) }{ \cos(x) }\]

OpenStudy (imstuck):

Now let's find a common denominator between the two:

OpenStudy (anonymous):



OpenStudy (imstuck):

I got something different from you...let me check mine and look at yours as well, ok?

OpenStudy (anonymous):

mine is probably wrong but alright, thanks.

OpenStudy (imstuck):

Don't say that...we all make oversights of important stuff...I did it twice today!

OpenStudy (anonymous):

lol thanks

OpenStudy (imstuck):

I got the same thing both times I checked so here's how I did it and see if you can follow my step by step, ok?

OpenStudy (anonymous):

alright

OpenStudy (imstuck):

We already established that csc = 1/sin and tan = sin/cos so I will go straight to the next part:

OpenStudy (imstuck):

\[\sin(x)\cos(x)(\frac{ 1 }{ \sin(x) }+\frac{ \sin(x) }{ \cos(x) })\]

OpenStudy (imstuck):

We will distribute the sin(x)cos(x) into the parenthesis to get this:

OpenStudy (imstuck):

...\[\frac{ \sin(x)\cos(x) }{ \sin(x)}+\frac{ \sin ^{2}(x)\cos(x) }{ \cos(x) }\]

OpenStudy (anonymous):

ok got that part

OpenStudy (imstuck):

Now we need to put those over a common denominator. Do this by multiplying the first part of the expression by cos(x)/cos(x) and the second part by sin(x)/sin(x):

OpenStudy (anonymous):

fingers crossed, hope i did it right :D

OpenStudy (imstuck):